Calculate the missing value. Report pH to 2 decimal places. Report [HCO3-] to 1 decimal place....

  1. Calculate the missing value. Report pH to 2 decimal places. Report [HCO3-] to 1 decimal place.
  2. Indicate whether the patient has an acidosis or an alkalosis.
  3. Indicate whether the primary disturbance is respiratory or metabolic and whether it is uncompensated, partially compensated, or fully compensated.

This patient hasn’t eaten in over 24 hours.

pO2              =        110 mm Hg

pCO2           =        30 mm Hg

[HCO3]       =        17 mmole/L

base excess =        -2 mmole/L

pH =

acid/base status:
Answer #1

pH = pK + log [HCO3]/ [PCO2]

pK = 6.1

HCO3 = 17mmhg

PCO2= 30mmhg

pH = 6.1 + log 17/ 30 , = 6.1 - 0.251

= 5.8, case of metabolic acidosis and is partially compensated